# Early-stopping for Transformer model training

### v1

@@ Appendix C Dynamic Evolution of Spectral Parameters and Training Phase Division
C.1 Workflow for Spectral Parameter Analysis of Attention Layer Matrices
On a log-log plot, a power-law distribution appears as a straight line, where the parameter $\alpha$ corresponds the negative slope of this line (Clauset et al., 2009; Klaus et al., 2011; Alstott et al., 2014; Martin and Mahoney, 2021; Yang et al., 2022). The value of $\alpha$ directly characterizes the distribution’s “tail heaviness”: a larger $\alpha$ implies a steeper slope and thus a faster decay of the distribution. To ensure a comprehensive analysis, we initially performed PL fitting on the Q, K, and V matrices from all attention mechanisms—encoder self-attention, decoder self-attention, and cross-attention. After a preliminary evaluation, we found that the spectral parameter evolution of the en.0.s.a.v matrix was the clearest and most consistent, whereas the Q and K matrices exhibited higher variability and layer-specific fluctuations. Based on this observation, and to further probe the essential training dynamics, we selected the spectral parameters of the V matrix as the primary indicators for subsequent analysis and for constructing the early-stopping mechanism. This choice aligns with the representative matrix analysis presented in Appendix B.
Our analysis methodology is as follows:
@@ C.1 Workflow for Spectral Parameter Analysis of Attention Layer Matrices
    Data Collection: During the training of the two experimental setups (T1_Ch-En, and T2_Ch-En), we periodically extracted the Q, K and V weight matrices from all attention mechanisms at checkpoints at multiple training epochs. Specifically, matrices were extracted every 50 epochs, resulting in six sampled epochs per experiment.
@@
C.2 Visual Analysis of the Evolution Process of Spectral Parameters
To perform a comprehensive multi-dimensional analysis of the dynamic evolution of the Power-Law fitting parameters ($\alpha$ and $x_{\min}$), we developed a set of four complementary visualization techniques. For each of the two experimental configurations, these methods are presented in subplots (a)–(d) of the corresponding result figures. The objective is to capture the macroscopic trends, local stability, and fine-grained fluctuations of the parameters throughout training. The visualization results for the two experimental setups are shown in Figures 8, 9, 10, and 11.
  • Raw Dynamic Observation (Subplots (a) & (c) - Raw Data Plots): Subplots (a) and (c) directly visualize the unprocessed fitted parameters as scatter plots connected by lines over the training epochs. This provides the most immediate view of parameter fluctuations across the entire training process.
  • Macroscopic Trend Analysis (Subplot (b) - Mean Value Evolution Plot): To capture the global convergence trends, we compute the mean values of the parameters over predefined epoch groups (e.g., epochs 1-6, epochs 50-55). Subplot (b) then plots these mean values as a function of training progress, thereby illustrating the long-term trajectory of parameter evolution.
  • Local Stability Analysis (Subplot (d) - Standard Deviation Evolution Plot): To assess the degree of local stability, we calculate the standard deviation of the parameters within each sampling group. Subplot (d) depicts these values over time, where larger deviations indicate strong fluctuations and instability, while smaller deviations suggest convergence and stabilization of the parameters.
@@ C.2 Visual Analysis of the Evolution Process of Spectral Parameters
  • Shallow layers are more stable than deep layers, and the cross-attention mechanism exhibits the most significant fluctuations: Across both the T1 and T2 models, the $\alpha$ trajectories of shallower layers (e.g., Layer-0) are notably smoother and more stable than those of deeper layers (e.g., Layer-1 and Layer-2) (see Figures 8 and 10). In contrast, the $\alpha$ values of the decoder’s cross-attention layers (Decoder-cross) display the strongest randomness and volatility throughout training.This behavior likely reflects the inherently dynamic alignment process between the source and target language representations performed by the cross-attention mechanism, leading to slower convergence of the spectral properties of its weight matrices.

### v2

@@ Appendix C Dynamic Evolution of Spectral Parameters and Training Phase Division
This section analyzes the dynamic evolution of spectral parameters ($\alpha$ and $x_{\min}$) in Transformer weight matrices to uncover the intrinsic dynamics across different training phases. By establishing a comprehensive workflow involving data extraction, SVD decomposition, and Power-Law fitting, combined with multi-dimensional visualization techniques, we systematically evaluate the stability of various attention mechanisms. Our analysis reveals three key findings: (1) shallow layers are significantly more stable than deep layers; (2) the cross-attention mechanism exhibits the most pronounced fluctuations; and (3) the shallow encoder $V$ matrix (en.0.s.a.v) demonstrates the most consistent convergence trend and a unique heavy-tailed evolution pattern.(This result aligns with the representative matrix analysis presented in Appendix B.1.)
On a log-log plot, a Power-Law distribution appears as a straight line, where the parameter $\alpha$ corresponds to the negative slope. The value of $\alpha$ directly characterizes the distribution’s “tail heaviness”: a larger $\alpha$ implies a steeper slope and a faster decay. To capture training dynamics, we followed this workflow:
@@
    Data Collection: During the training of the two experimental setups (T1_En-Ch, and T2_En-Ch), we periodically extracted the Q, K and V weight matrices from all attention mechanisms at checkpoints at multiple training epochs. Specifically, matrices were extracted every 50 epochs, resulting in six sampled epochs per experiment.
@@
To perform a comprehensive multi-dimensional analysis of the dynamic evolution of the Power-Law fitting parameters ($\alpha$ and $x_{\min}$), we developed a set of four complementary visualization techniques. For each of the two experimental configurations, these methods are presented in subplots (a)–(d) of the corresponding result figures. The objective is to capture the macroscopic trends, local stability, and fine-grained fluctuations of the parameters throughout training. The visualization results for the two experimental setups are shown in Figures 9, 10, 11, and 12.
  • Raw Dynamic Observation (Subplots a & c): Directly visualizes the unprocessed fitted parameters over training epochs to show immediate fluctuations.
  • Macroscopic Trend Analysis (Subplot b): Plots the mean values of parameters over predefined epoch groups, illustrating the long-term trajectory of parameter evolution.
  • Local Stability Analysis (Subplot d): Depicts the standard deviation of parameters within sampling groups; larger deviations indicate instability, while smaller ones suggest convergence.
@@
  • Shallow layers are more stable than deep layers, and the cross-attention mechanism exhibits the most significant fluctuations: Across both the T1 and T2 models, the $\alpha$ trajectories of shallower layers (e.g., Layer-0) are notably smoother and more stable than those of deeper layers (e.g., Layer-1 and Layer-2) (see Figures 9 and 11).In contrast, the $\alpha$ values of the decoder’s cross-attention layers (Decoder-cross) display the strongest randomness and volatility throughout training.This behavior likely reflects the inherently dynamic alignment process between the source and target language representations performed by the cross-attention mechanism, leading to slower convergence of the spectral properties of its weight matrices.
@@
Appendix D Technical Proofs for recognition criteria for Heavy-Tail based on PL fitting
D.1 Proof of Lemma 3.1
$P(\sup_{x\in\mathbb{R}}|\hat{F}_{n}(x)-F(x)|>\epsilon)\leq 2e^{-2n\epsilon^{2}}.$ (D.1)
Consequently, we obtain
$P(d>\epsilon)\leq 2e^{-2n_{\text{tail}}\epsilon^{2}}.$
D.2 Proof of Proposition 3.1
Consider the KS distance
$=d+\sup_{x\geq x_{min}}|F(x;\alpha,x_{min})-F(x;\hat{\alpha},x_{min})|.$ (D.2)
Log-likelihood function is
[…]
Combining these results, we obtain that
$=$ $\left(\frac{x}{x_{min}}\right)^{-\alpha+1}\left(\left(\frac{x}{x_{min}}\right)^{-(\hat{\alpha}-\alpha)}-1\right),$ (D.3)
$F(x;\alpha,x_{min})-F(x;\hat{\alpha},x_{min})=O_{p}\left(\frac{1}{\sqrt{n_{tail}}}\right).$ (D.4)
Hence, the convergence rate of the distance $\hat{d}$ is
$\hat{d}=O_{p}\left(\frac{1}{\sqrt{n_{\text{tail}}}}\right).$
Appendix E Determination of Threshold C via Monte Carlo Simulation
Step 1: Set Simulation Parameters
  • Number of Monte Carlo Iterations (k): We set $k=10,000$ repetitions.
Step 2: Generate Synthetic Data
$x=x_{\min}\cdot(1-u)^{-\frac{1}{\alpha_{0}-1}},$
Step 3: Calculate the Normalized Statistic S
  1. 1.
  2. 2.
  3. 3.
    Calculate the normalized KS statistic S:
Step 4: Select Threshold C
Following the steps outlined above, the results of our simulation are presented in Figure 13. The simulation results indicate that, across 10,000 Monte Carlo iterations for various values of $\alpha$ and tail data sizes ($n_{\text{tail}}$), the upper bound of the normalized KS statistic remains highly stable and does not exceed 2. Consequently, we select C=2 as the critical threshold.
Figure 13: Monte Carlo simulation for calibrating the threshold constant C. Each subplot displays the distribution of the statistic $S=\hat{d}\cdot\sqrt{n_{\text{tail}}}$ over 10,000 runs, for different true values of $\alpha$ and tail sizes ($n_{\text{tail}}$). The vertical blue line indicates the proposed threshold.
Appendix F Evolutionary Trajectory of the Early-Stopping Metric for T2_En-Ch 9.62M and T3_En-Ch 19.20M
Figure 14: Evolutionary Trajectory of the Early-Stopping Metric. (a) and (b) describe the variation of d and d* with training epochs for T2_En-Ch 9.62M and T3_En-Ch 19.20M, respectively; while (b) and (d), illustrate the changes in the difference between d* and d across training epochs for these two models, respectively.